OpenStudy (anonymous):

What is the area of a regular nonagon with a radius of 14 in.? I will include my work below. I just need to know where I am going wrong.

OpenStudy (anonymous):

Okay so I took 360/9 and got 40 and divide that by 20 to get the degree that I will work with. Then I used the equation sin20=x/14 and got 4.788282007. Then I used cos20=x/14=13.15569669. So then I would have to use A=1/2ap. and i get 31.49659288 but that's wrong.